OpenStudy (anonymous):

What are the asypmtotes of the hyperbola (x-1)^(2)/(25)-(y-4)^(2)/(16)= ?

OpenStudy (anonymous):

In general the asymptotes of a hyperbola are y = (b/a) and y = - (b/a).

OpenStudy (anonymous):

sorry they are y=+/-(b/a)x

OpenStudy (anonymous):

5y=24-4x and 5y=4x+16

OpenStudy (anonymous):

Where did you come up with that equation?

OpenStudy (anonymous):

y=(b/a)x and y=-(b/a)x are the asymptotes the the "general" hyperbola

OpenStudy (anonymous):

but where did you pull the numbers from?

OpenStudy (anonymous):

y-4=+/-(4/5)(x-1)

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

yes.